# Skrull: Towards Efficient Long Context Fine-tuning through Dynamic Data Scheduling

### v1

@@ 5 Evaluation
We conduct experiments using a testbed consisting of 4 nodes interconnected via a high-performance InfiniBand network, with each node equipped with 8 Nvidia H100 GPUs connected via 900GB/s NVLink. Then, We implement Skrull on top of DeepSpeed, a state-of-the-art distributed LLM training system and enable Zero-2 optimization as our baseline. We evaluation our optimizations on Qwen2.5-0.5B and Qwen2.5-7B using the three real-world datasets described in Section 3.1. Although Wikipedia and LMsysChat1M are not specifically gathered for Long-SFT, we still choose them as our evaluation datasets due to their long-tail distribution, which is exactly identical to Meta’s in-house Long-SFT dataset [10]. In contrast, ChatQA2-long-SFT dataset [22] is specifically gathered for Long-SFT and exhibits bimodal distribution of data length, which is also similar to the dataset mentioned in [19]. Through offline profiling, we configure the BucketSize to 26K and 13K for Qwen2.5-0.5B and Qwen2.5-7B, respectively. Further details regarding BucketSize configuration can be found in Appendix A.1. All the experiments share the same training settings with <DP=4, CP=8, BatchSize=64>, zero-2 enabled and selective recomputation strategy except for training Qwen-2.5-7B with ChatQA2-long-SFT dataset. Due to the increased memory requirements, we adjust its parallel settings with <DP=2, CP=16, BatchSize=40>.
@@
Figure 3 illustrates the speedup achieved by Skrull, with the performance measured in terms of average iteration time. Experimental results demonstrate that Skrull outperforms DeepSpeed by an average of 3.76x and achieves peak improvement of 7.54×. The average speedups for Qwen-0.5B and Qwen-7B are 5.50x and 2.03x, respectively. We attribute this difference to the variation in BucketSize, which directly influences the valid data scheduling space. Additionally, from the perspective of datasets, the performance on Wikipedia and LMsysChat1M are similar due to the similar data distribution, which both exhibit long-tail feature. In this distribution, the short sequences dominate the datasets thus showcasing more optimization potential. In contrast, the long sequences also account for the majority in ChatQA2-Long-SFT dataset, which exhibits bimodal distribution, leading to relatively small optimization space. Specifically, when training Qwen-7B with this datasets, the major sequence length exceeds the BucketSize thus leading to limited speedup. We can further extend the BucketSize by combining more optimization techniques like parameter-efficient fine-tuning (PEFT), which we leave to future works.
@@
Performance Impact of BatchSize.
Figure 4: Speedup vs. batch size.
To investigate the performance impact of batch size, we conduct experiments on ChatQA2-long-SFT using Qwen2.5-0.5B. As shown in Figure 4, as the batch size increases from 8 to 54, the end-to-end speedup also improves. We attribute this performance gain to the expanded scheduling scope achieved with larger batch sizes. However, as the batch size continues to increase, the sampled batches gradually converge to the sequence length distribution of the dataset, causing the performance gains stabilized within a reasonable range.
@@ 6 Related Works
A lot of researches are conducted to the efficient long-SFT. However, most works are focus on the perspective of data engineering [24, 3, 19, 10]. Another type of works are parameter efficient finetuning (PEFT) [11, 5] and Skrull is also effective for this methods. From the perspective of training system, LongAlign [3] adopts a sorted batching to optimize system efficient while break equivalence during training. Chunkflow [23] organize sequences into fixed size chunks, enabling controllable peak memory consumption and reduced pipeline bubbles. Additionally, some works employ dynamic parallelism settings [8] to handle varying length sequences, which is similar to long-SFT. In contrast, Skrull adopts fixed parallelism settings and is orthogonal to those methods.
@@ 7 Conclusion
In this paper, we provide a new prospective of data scheduling to enhance the training efficiency in Long-SFT scenarios. The heterogeneous data distribution in Long-SFT poses dilemmas for existing training systems on configuring parallelism strategies and ensuring the load balance. To tackle those challenges, we propose Skrull, a dynamic data scheduler dedicated for Long-SFT. Through dynamic data scheduling, Skrull achieves efficient training on both long sequences and short sequences. Additionally, we formulate the scheduling process as a joint optimization and adopt a lightwight scheduling algorithm. Experimental results demonstrate that Skrull outperforms DeepSpeed by 3.76x on average (up to 7.54x) in real-world long-SFT. Furthermore, we believe that Skrull can serve as an effective solution in other scenarios especially when dealing with mixture of long and short training data, such as reinforcement learning from human feedback (RLHF).
Appendix A Performance Modeling
A.1 Memory Estimation
@@
We first analyze the memory consumption during LLMs training. The memory consumption can be roughly categorized into two components: the static memory and the dynamic memory. The static memory, which typically includes model parameters and optimizer states, remains roughly constant throughout the training process given specific model configurations and parallelism strategies. In contrast, the dynamic memory or activation memory, varies with the input workload. In transformer architectures, activation memory is proportional to the sequence length. For instance, the Linear module, LayerNorm and Attention module (using FlashAttention [7, 6]) exhibit a linear relationship with sequence length. Therefore, we can estimate activation memory for a given sequence length $S$ using the following equation:
@@ A.3 Communication Estimation
Appendix B Details of Heuristic Algorithm
Due to the limited page, We list the omitted function definition in heuristic scheduling algorithm of CP in Algorithm 3.
Algorithm 3 Function Definations in scheduling algorithm for DACP
1:SeqNum $K$, SeqLens $S[K]$, Buckets $C$, CP degree $N$, Loads $L[N]$, RemainBucket $RB[N]$, DACP scheduling result $ret$
2:function UpdateLocal($idx$, $rank$)
3:     $RB[rank]\leftarrow RB[rank]-S[idx]$triangleright Update remaining bucket capacity
4:     $L[rank]\leftarrow L[rank]+FLOPs(S[idx])$triangleright Update current load
5:end function
6:function UpdateAll($idx$)
7:     for $i=0$ to $N-1$ do
8:         $RB[i]\leftarrow RB[i]-S[idx]/N$triangleright Distribute across all buckets
9:         $L[i]\leftarrow L[i]+FLOPs(S[idx],N)$triangleright Update all loads
10:     end for
11:end function
12:function RollBack($rank$)
13:     for $i=0$ to $K-1$ do
14:         if $ret[i]==rank$ then
15:              $ret[i]\leftarrow-1$triangleright Distribute the sequence
16:              $RB[rank]\leftarrow RB[rank]-S[i]+S[i]/N$
17:              $L[rank]\leftarrow L[rank]-FLOPs(S[i])+FLOPs(S[i],N)$
18:              return $True$triangleright Success Roll-back
19:         end if
20:     end for
21:     return $False$triangleright Roll-back Failed
22:end function

### v2

@@ 5 Evaluation
We conduct experiments using a testbed consisting of 4 nodes interconnected via a high-performance InfiniBand network, with each node equipped with 8 Nvidia H100 GPUs connected via 900GB/s NVLink. Then, We implement Skrull on top of DeepSpeed, a state-of-the-art distributed LLM training system and enable Zero-2 optimization as our baseline. Additionally, we implement sorted batching method in LongAlign [bai_longalign_2024] for more comparison, which sort the dataset by sequence length and select random consecutive groups for each batch to improve the long-SFT training efficiency. We evaluation our optimizations on Qwen2.5-0.5B and Qwen2.5-7B using the three real-world datasets described in Section 3.1. Although Wikipedia and LMsysChat1M are not specifically gathered for Long-SFT, we still choose them as our evaluation datasets due to their long-tail distribution, which is exactly identical to Meta’s in-house Long-SFT dataset [llama3_tech_report_2024]. In contrast, ChatQA2-long-SFT dataset [dataset_chatqa2-long-sft-data_2025] is specifically gathered for Long-SFT and exhibits bimodal distribution of data length, which is also similar to the dataset mentioned in [qwen25_tech_report_2024]. Through offline profiling, we configure the BucketSize to 26K and 13K for Qwen2.5-0.5B and Qwen2.5-7B, respectively. Further details regarding BucketSize configuration can be found in Appendix C.1. All the experiments share the same training settings with <DP=4, CP=8, BatchSize=64>, zero-2 enabled and selective recomputation strategy except for training Qwen-2.5-7B with ChatQA2-long-SFT dataset. Due to the increased memory requirements, we adjust its parallel settings with <DP=2, CP=16, BatchSize=40>. The global batch size is equal to DP size multiplied by BatchSize. Due to the limited page, we list precision validation in Appendix A.
@@
Figure 3 illustrates the speedup achieved by Skrull, with the performance measured in terms of average iteration time. Skrull significantly outperforms the sorted batching strategy Experimental results demonstrate that Skrull outperforms DeepSpeed and sorted batching method by an average of 3.76x and achieve peak improvement of 7.54×. The average speedups for Qwen-0.5B and Qwen-7B are 5.50x and 2.03x, respectively. We attribute this difference to the variation in BucketSize, which directly influences the valid data scheduling space. Also, Skrull outperforms sorted batching method by an averge of 3.45x with the peak improvement of 6.85x. Additionally, from the perspective of datasets, the performance on Wikipedia and LMsysChat1M are similar due to the similar data distribution, which both exhibit long-tail feature. In this distribution, the short sequences dominate the datasets thus showcasing more optimization potential. In contrast, the long sequences also account for the majority in ChatQA2-Long-SFT dataset, which exhibits bimodal distribution, leading to relatively small optimization space. Specifically, when training Qwen-7B with this datasets, the major sequence length exceeds the BucketSize thus leading to limited speedup. We can further extend the BucketSize by combining more optimization techniques like parameter-efficient fine-tuning (PEFT) [hu_lora_2022, chen_longlora_2024].
@@
Performance Impact of BatchSize and BucketSize.
Figure 4: Left (a) shows the performance impact under different BatchSize and BucketSize. Right (b) shows Skrull’s effectiveness and compatibility with Lora in larger models.
To investigate the performance impact of BatchSize and BucketSize, we conduct experiments on ChatQA2-long-SFT using Qwen2.5-0.5B with the default setting of <DP=1, CP=8, BatchSize=64, BucketSize=26K>. As shown in Figure 4(a), we adjust the BatchSize from 8 to 54 and the end-to-end speedup also improves. We attribute this performance gain to the expanded scheduling scope afforded by larger batch sizes. However, as the BatchSize increases further, the sampled batches gradually converge to the sequence length distribution of the dataset, causing the performance gains stabilized within a reasonable range. Additionally, we also evaluate the effect of BucketSize. Figure 4(a) shows that increasing the BucketSize from 8K to 32K (values in parentheses) improves the speedup until an out-of-memory (OOM) error occurs. This indicates that while a larger BucketSize enhances performance, it also raises the risk of OOM errors. Therefore, it is important to set appropriate BucketSize, highlighting the importance of performance modeling module in Skrull.
Case Study
Table 3: Scheduling Strategies Comparison
Method Speedup
[…]
RR w/o roll-back OOM
[…]
Skrull w/o roll-back OOM
Table 4: Latency Comparison per Iteration
Iteration Baseline $R$ Skrull
1 (36, 36) (10, 49) (21, 51)
2 (40, 40) (29, 47) (37, 50)
3 (35, 35) (14, 45) (32, 49)
4 (46, 46) (46, 47) (45, 49)
In this section, we present a quantitative analysis of the training process of Skrull. We conduct experiments using the Qwen2.5-0.5B model with the ChatQA2-Long-SFT dataset under the configuration <DP=1, CP=8, BatchSize=64>. To evaluate the effectiveness of our heuristics implementation, we compare it against a round-robin (RR) scheduling strategy, which assigns sequences in a simple round-robin manner (details in Appendix B.2). Additionally, we test both scheduling algorithms with and without roll-back mechanism to further assess the trade-off design in Skrull. As shown in Table 4, Skrull significantly outperforms the RR scheduling. We analyze this result by examining quantitative data in the first four iterations. Table 4 (presented in tuple format) reports the minimum and maximum peak memory usage (in gigabytes) across all GPUs during each iteration. Compared to the baseline, both Skrull and RR scheduling allocate more sequences locally at the cost of increased memory imbalance. While such imbalance is acceptable as long as it dose not exceed memory capacity, it raises the risk of OOMs. Therefore, as shown in Table 4, without the roll-back mechanism, both scheduling strategies result in OOMs, underscoring the importance of this safeguard. In contrast to RR scheduling, Skrull achieves a better computational balance (indicated by the speedup) while adhering to memory constraints, demonstrating the effectiveness of its trade-off design in Skrull.
@@ 6 Related Works
From the perspective of data engineering, those works [zhao_longskywork_2024, bai_longalign_2024, qwen25_tech_report_2024, llama3_tech_report_2024] involve meticulously gathering training datasets for long context fine-tuning. From the perspective of training system, LongAlign [bai_longalign_2024] adopts a sorted batching strategy to optimize system efficiency in long context fine-tuning phase. Chunkflow [chunkflow] organize the training data into fixed size chunks, enabling controllable peak memory consumption and reduced pipeline bubbles. Additionally, some works employ dynamic parallelism settings [hotspa_sosp24] to handle varying length sequences, which is similar to long-SFT. In contrast, Skrull adopts fixed parallelism settings and is orthogonal to those methods. Another type of works are parameter efficient finetuning (PEFT) [hu_lora_2022, chen_longlora_2024] and Skrull is also effective for this methods.
@@ 7 Conclusion
In this paper, we provide a new perspective of data scheduling to enhance the training efficiency in Long-SFT scenarios. The heterogeneous data distribution in Long-SFT poses dilemmas for existing training systems on configuring parallelism strategies and ensuring the load balance. To tackle those challenges, we propose Skrull, a dynamic data scheduler dedicated for Long-SFT. Through dynamic data scheduling, Skrull achieves efficient training on both long sequences and short sequences. Additionally, we formulate the scheduling process as a joint optimization and adopt a lightwight scheduling algorithm. Experimental results demonstrate that Skrull outperforms DeepSpeed by 3.76x on average (up to 7.54x) in real-world long-SFT. Furthermore, we believe that Skrull can serve as an effective solution in other scenarios especially when dealing with mixture of long and short training data, such as reinforcement learning from human feedback (RLHF).
Acknowledgments
This work is supported by the following funding: Beijing Natural Science Foundation (4254087), National Science Foundation of China (62502501, 92270206, 62372435), China National Postdoctoral Program for Innovative Talents (BX20240383), Strategic Priority Research Program of Chinese Academy of Sciences (XDB0500102), and Alibaba Research Intern Program. The model training were performed on the robotic AI-Scientist platform of Chinese Academy of Science and Alibaba Cloud Platform for AI (PAI).
Appendix A Precision Validation
To evaluate the loss equivalence, we compare the loss carve between Skrull and standard training method when training Qwen2.5-0.5B on LMsysChat1M. The data scheduling in Skrull alter the accumulation order and we can observe slightly numerical differences due to the non-associativity of floating-point operations. However, Skrull do not alter any contents and orders in each global batch, the optimization trajectory remains equivalent. Therefore, as shown in Figure 5, Skrull does not influence the convergence.
Figure 5: Loss comparison between Skrull and standard training method.
Appendix B Heuristic Scheduling Algorithm
B.1 Function Definition of Heuristic Algorithm
Algorithm 3 Function Definations in scheduling algorithm for DACP
1:SeqNum $K$, SeqLens $S[K]$, Buckets $C$, CP degree $N$, Loads $L[N]$, RemainBucket $RB[N]$, DACP scheduling result retret
2:function UpdateLocal(idxidx, rankrank)
3:  $RB[rank]\leftarrow RB[rank]-S[idx]$ $\triangleright$ Update remaining bucket capacity
4:  $L[rank]\leftarrow L[rank]+FLOPs(S[idx])$ $\triangleright$ Update current load
5:end function
6:function UpdateAll(idxidx)
7:  for $i=0$ to N1N-1 do
8:   $RB[i]\leftarrow RB[i]-S[idx]/N$ $\triangleright$ Distribute across all buckets
9:   $L[i]\leftarrow L[i]+FLOPs(S[idx],N)$ $\triangleright$ Update all loads
10:  end for
11:end function
12:function RollBack(rankrank,RBRB,$L$)
13:  for $i=0$ to K1K-1 do
14:   if $ret[i]==rank$ then
15:     $ret[i]\leftarrow-1$ $\triangleright$ Distribute the sequence
16:     $RB[rank]\leftarrow RB[rank]-S[i]+S[i]/N$
17:     if $L$ is not None then
18:      $L[rank]\leftarrow L[rank]-FLOPs(S[i])+FLOPs(S[i],N)$
19:     end if
20:     return TrueTrue $\triangleright$ Success Roll-back
21:   end if
22:  end for
23:  return FalseFalse $\triangleright$ Roll-back Failed
24:end function
B.2 Round-robin Scheduling Algorithm
Algorithm 4 Round-Robin Scheduling Algorithm
1:Input: SeqNum NN, SeqLens $S$, Buckets $C$, WorldSize wsws
2:Output: PartitionIds $P$
3:for $i=1$ to m1m-1 do
4:  $t\leftarrow\text{FindMaxBucketsIds()}$
5:  if $C[t]\geq S[i]$ then
6:   $P[i]\leftarrow t$ $\triangleleft$ fit the max bucket
7:  else
8:   $j\leftarrow\text{FindMinBucketsIds()}$
9:   if $C[j]\geq S[i]/ws$ then
10:     $P[i]\leftarrow-1$ $\triangleleft$ partition the sequence
11:   else
12:     Assert RollBack(j,C)(j,C) $\triangleleft$ with roll-back
13:     $i\leftarrow i-1$
14:     continue
15:   end if
16:  end if
17:end for
Appendix C Performance Modeling
C.1 Memory Estimation
@@
We first analyze the memory consumption during LLMs training. The memory consumption can be roughly categorized into two components: the static memory and the dynamic memory. The static memory, which typically includes model parameters and optimizer states, remains roughly constant throughout the training process given specific model configurations and parallelism strategies. In contrast, the dynamic memory or activation memory, varies with the input workload. In transformer architectures, activation memory is proportional to the sequence length. For instance, the Linear module, LayerNorm and Attention module (using FlashAttention [flashattention_v1_2022, flashattention_v2_2024]) exhibit a linear relationship with sequence length. Therefore, we can estimate activation memory for a given sequence length $S$ using the following equation: